OpenStudy (anonymous):

using the concept of slope to find "t" such that the three points are collinear; (1,-4), (t,3), (5,10)

OpenStudy (amistre64):

make the slope from the outer points match with what the middle needs to be

OpenStudy (amistre64):

in other words, slope the outers and use the middle to adjust after you determine a suitable equation for the line

OpenStudy (anonymous):

i found the slop of the outers its 7/2. however my dilema is i dont know which equation to use to find "t"

OpenStudy (anonymous):

slope*

OpenStudy (amistre64):

7/2 is good; now we would devise an equation that fits all the point onto it. since we know the slope, and have alot of given points; lets use one of the points to form the POINT SLOPE form.

OpenStudy (amistre64):

y - Py = m( x - Px); ill use (1,-4) y - -4 = 7/2(x - 1) 2y + 8 = 7 (x - 1) (2y + 8)/7 = x-1 (2y + 8)/7 +1 = x = t now when y=3 we can determine a value for t (2(3) + 8)/7 +1 = t (6+8)/7 +1 = t 14/7 +1 = t = 3

OpenStudy (amistre64):